MONT 107Q – Thinking About Mathematics First Writing Assignment

The firstt of the larger paper assignments this semester will be devoted to one of the two topics sketched below, or to a topic of your choice. Your essay should be at least 4 pages in length (double-spaced, single-sided). Please submit via Google Docs or as a MS Word .doc or .docx file to jlittle@holycross.edu by 5:00pm on Monday, February 6.

Paper topic 1

In All the Pretty Horses, do you think we have to agree that the main character, John Grady Cole, grows up? Why or why not? It is certainly the case that he goes to Mexico with dreams about what it means to be a man and a cowboy. What happens to those dreams as a result of the trip? Does he change? If so, what has he changed into by the end of the story? Would it be right to say he is a hero? What would it mean to call him a hero in the terms of the society and culture(s) portrayed in the novel?

Paper topic 2

Some recent literary criticism has argued that “Western” films and novels deal primarily with masculine issues and leave the concerns and experiences of women almost entirely out of the picture. Women appear only as plot devices or as sexual objects and the main focus is on the importance of masculinity, standards of male conduct, and the problem of creating an authentic male identity. (And if women do take a key part, the underlying “point” is how they are taking on masculine roles – think of the role of “Emma Cullen” in last year’s The Magnificent Seven played by Haley Bennett if you saw that film.) To what extent do you think this is true about All the Pretty Horses? Do women have important roles? If so, who are the women, what are their roles, and how are they different from male roles? What sorts of relationships does John Grady Cole have with women? What are his attitudes about women?

Paper topic 3

Identify some aspect of All the Pretty Horses that you find especially interesting or noteworthy, frame a thesis in response to a question about that aspect, and present a argument supporting your proposed answer. I would be happy to discuss possible ideas before you begin to write.

Guidelines for this assignment

Follow these specifications:

1. It is OK to discuss preliminary ideas for what you want to say with your classmates or with me if you want. I will also have time to read preliminary drafts and get feedback to you again on this paper.

2. As usual, start by clearly stating the main idea you want to get across in your paper – its thesis. It should be clear to the reader which of the two topics you are addressing.

3. In the following paragraphs, by drawing on our class discussions, and your reading of All the Pretty Horses, present your evidence for your conclusions.

4. You may present evidence consisting of quotations of passages from the book, together with interpretations of those passages that come from your own thinking about them. Your evidence should be as specific, detailed, correct, and as varied as possible (focus on small details and mention the little things that support your point of view). If you do use quotations, identify them sufficiently carefully that I can easily track them down. (For quotations from All the Pretty Horses, a page number in the Vintage paperback edition is sufficient.)

5. Your essay should be well-organized as a whole and “flow” from one point to the next. Following an overall outline laid out in advance, subdivide your paper into paragraphs that each have one main point. Do not repeat pieces of evidence unless there is a very good reason for doing so. Do not get ahead of yourself. Say one thing at a time. If you decide you want to add something, do not just tack it on somewhere. Instead, figure out where it fits in the overall scheme of your argument and put it where it belongs.

6. As always, your sentences should be clear, readable, and not overly wordy. Aim for a reasonably formal but lively style. For this assignment, use of the first person is more or less expected while you are presenting your opinion about the effectiveness of the pairing of Othello with Othello in the Seraglio. Generally, colloquial ways of saying things that are acceptable in casual conversation are not as appropriate in this kind of writing, but if you have a really good reason for using a colloquial expression and it fits well, that’s OK. Also, please do not use a thesaurus to find impressive-sounding words when simpler alternatives are available. An important lesson to learn for writing in college is that thesauri do not capture the nuances in meanings that underlie good writing. If you must use a thesaurus, also look up any words you select in a dictionary to make sure they mean exactly what you think they do!

7. And as on previous assignments, errors in mechanics and usage should not annoy the reader. (As a warning, I tend to be easily annoyed by this sort of thing!) Try as much as possible to avoid mistakes in grammar, spelling, punctuation, and word choice.